A 12-ounce soft drink can has volume 21.66 cubic inches. If the top and botton are twice as thick as the sides, find the dimensions (radius and height) that minimize the ammount of metal used in the can.
1
Expert's answer
2012-07-06T07:19:23-0400
V = Pi · r^2 · h
h = V / (Pi · r^2)
m = a · (Ss + 2 · St + 2 · Sb) = a ·
(2 · Pi · r · h + 2 · Pi · r^2 +
2 · Pi · r^2) = a · (2 · Pi · r · V / (Pi ·
r^2) + 4 · Pi · r^2) = a ·
(2 · V / r + 4 · Pi · r^2)
Finding the minimum
of 2 · V / r + 4 · Pi · r^2:
2 · V / r + 4 · Pi · r^2 = 0
V + 2 · Pi · r^3
= 0
r = cubrt(V / (2 · Pi)) = cubrt(21.66 / (2 · 3.14)) = cubrt(21.66 /
(2
· 3.14)) = 1.51 inches
h = V / (Pi · r^2) = 21.66 / (3.14 · 1.51 ·
1.51) = 3.03 inches
Answer: 1.5 inches, 3 inches.